OpenStudy (anonymous):

Solve 5^x + 5 = 9^x a.) x ≈ -3.63 b.) x ≈ -0.28 c.) x ≈ 0.07 d.) x ≈ 13.69

OpenStudy (anonymous):

a.) x ≈ -3.63 b.) x ≈ -0.28 c.) x ≈ 0.07 d.) x ≈ 13.69

Directrix (directrix):

I think this problem does not go with these options OR these options do not go with the posted problem. Please check. Thanks.

OpenStudy (anonymous):

no its correct

Directrix (directrix):

I think you added the exponents. :) •Edited 2 minutes ago •

OpenStudy (anonymous):

yeah sorry about that

OpenStudy (anonymous):

The answer to \[5^{5+x}=9^x \]is x=13.6907